### v1

@@
The requirements of unitarity and causality lead to significant constraints on the Wilson coefficients of an EFT expansion, known as positivity bounds. Their standard derivation relies on the crucial assumption of polynomial boundedness on the growth of scattering amplitudes in the complex energy plane, which is a property satisfied by local QFTs, and by weakly coupled string theory in the Regge regime. The scope of this work is to clarify the role of locality by deriving generalized positivity bounds under the assumption of exponential boundedness, typical of non-local QFTs where the Froissart-Martin bound is usually not satisfied. Using appropriately modified dispersion relations, we derive new constraints and find regions in the EFT parameter space that do not admit a local UV completion. Furthermore, we show that there exist EFTs that satisfy IR causality and at the same time can admit a non-local UV completion, provided that the energy scale of non-locality is of the same order or smaller than the EFT cutoff. Finally, we provide an explicit example of a non-perturbative amplitude that simultaneously satisfies the properties of exponential boundedness, unitarity and causality.
@@
1 Introduction
Scattering amplitudes are very powerful tools for connecting the ultraviolet (UV) aspects of a fundamental theory with its properties and observables at low energies. To constrain the parameter space of viable effective field theories (EFTs) it is often not necessary to know the precise form of the UV completion. In many cases it is sufficient to assume certain physical properties in the UV regime, which can be translated into mathematical requirements on the $S$-matrix, from which constraints on low-energy physics can then be derived [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] (see also  [54] for a recent review). For example, standard requirements that are usually assumed are those of Lorentz invariance, unitarity and causality. These properties are encoded in a set of precise mathematical conditions, such as positivity of the imaginary part of the partial waves and analyticity of the $2\rightarrow 2$ scattering amplitude in the complex energy plane [55, 56, 57, 58].
A further assumption that is typically made on the UV completion is that of locality, which states that the growth of the scattering amplitude is polynomially bounded in the complex energy plane. For local gapped theories it can be proven that the amplitude for fixed physical momentum transfer is polynomially bounded by $|s|^{2}$ in the whole complex plane; this is known as Froissart-Martin bound [^1] It is worth to mention that Froissart [59] and Martin [60] initially showed that the amplitude was bounded by $|s|$ in the forward limit $t=0.$ Subsequently, Jin and Martin [61] showed that the amplitude for generic physical momentum transfer $t\leq 0$ is bounded by $|s|^{2}.$ In any case, in this work we refer to the latter as Froissart-Martin bound. [59, 60, 61]. Although for theories with massless states no rigorous general proof of this bound exists, it is common to assume that the growth of the amplitude for fixed physical $t$ is bounded by $|s|^{2}$ in the whole complex plane [^2] This can actually be verified for graviton-mediated amplitudes in $d>4$, using the method of eikonal resummation [62, 63].:
@@ 1 Introduction
where $B$ is some subdominant factor that can also depend on $s$.
To clarify the role of locality, it is instructive to investigate what happens when we allow for a non-polynomial growth of the scattering amplitude, i.e., for violations of the Froissart-Martin bound. In particular, we are going to consider amplitudes that are exponentially bounded in the complex $s$ plane for fixed physical momentum transfer:
@@
where $\alpha$ and $g$ are two constant positive parameters, and $C$ a subdominant factor that can also depend on $s$.
@@
QFTs in which scattering amplitudes are characterized by this type of exponential behavior could be related to Jaffe’s classification [64, 65, 66, 67, 68]: if $g<1/4$ the corresponding QFT is said to be localizable, whereas if $g\geq 1/4$ the QFT is said to be non-localizable. In particular, when $g=0$ we recover the standard local QFT description with polynomially bounded amplitudes. The name non-localizability is related to the fact that the fields must be delocalized through smeared test functions in order to make the position-space correlators (such as the Wightman function) finite and well-defined. In what follows we use the words local (non-local) and localizable (non-localizable) interchangeably.
The aim of this paper is to understand what type of constraints can be derived on the Wilson coefficients of the low-energy EFT expansion if the UV completion is assumed to be non-localizable according to (2). As a case study, we consider a gapless scalar theory whose full $2\rightarrow 2$ scattering amplitude $A(s,t)$ vanishes at the point $(s,t)=(0,0)$ and that in the forward limit is bounded as $A(s,t=0)<s^{2}e^{-\alpha s^{2g}}$ in the entire complex plane. Implementing modified dispersion relations that are suitable for exponentially bounded amplitudes, we will show that the standard positivity bounds [3, 54] are generalized and become dependent on the parameters $\alpha$ and $g$ in Eq. (2). In particular, we focus on the case $g=1.$ We will find that the generalized positivity bounds not always overlap with the standard ones but, depending on the value of $\alpha,$ there can exist also regions in the EFT parameter space that do not admit a local UV completion. We further compare these positivity constraints with those arising from the low-energy causality requirement and find large regions in the EFT parameter space where the theory is both causal and unitary, but can be non-localizable. After presenting a general investigation of positivity and causality bounds, we will construct explicit examples of non-perturbative scattering amplitudes that satisfy analyticity, crossing symmetry and partial-wave unitarity, but are not polynomially bounded.
The work is organized as follows.
Sec. 2:
We introduce the method of dispersion relations suitably adapted to the case of amplitudes that are exponentially bounded in the complex $s$ plane. In particular, we consider the case of forward limit, i.e. zero-momentum transfer, and specialize the study to $g=1$.
Sec. 3:
We derive the generalized positivity bounds in the forward limit by constraining the coupling constants $g_{4},$ $g_{6}$ and $g_{8}$ in the EFT expansion. We determine the dependence of the result on the value of the non-local scale $\alpha$, and make a comparison with the local case $\alpha=0.$
Sec. 4:
We generalize the analysis to the case of non-zero momentum transfer and derive further generalized positivity bounds that also constrain the coupling $g_{3}$ in the EFT expansion. In this case, additional assumptions – such as crossing symmetry – are necessary in order to write down a well-defined dispersion relation.
Sec. 5:
We compare the regions in the space of EFT parameters constrained by the positivity bounds with the regions limited by the requirement of low-energy causality. We show that non-local UV completions can be compatible with the requirement of causality only if the energy scale of non-locality is of the same order or smaller than the EFT cutoff.
Sec. 6:
We comment on the implications of our results. In particular, we remark that our findings challenge the usually-claimed uniqueness of string theory as the only UV completion of gravity that is consistent with the requirements of unitarity and causality. Finally, we draw our conclusions.
App. A:
In this appendix we construct an explicit example of a non-perturbative amplitude that satisfies the properties of analyticity, causality, partial-wave unitarity, but violate polynomial boundedness.
App. B:
More mathematical details on the analytic derivation of the bound for the coupling $g_{3}$ are provided.
Conventions and notations.
@@
Throughout the entire paper, we work with the mostly plus metric signature ($+$)(-$+$)( - + + + ) and adopt the Natural units system $(\hbar=1=c).$ The central object of our investigation is the elastic $2\rightarrow 2$ scattering amplitude $A(s,t,u),$ where $s=-(p_{1}+p_{2})^{2},$ $t=-(p_{1}-p_{3})^{2},$ $u=-(p_{1}-p_{4})^{2}$ are the Mandelstam variables, $p_{1},$ $p_{2}$ are the ingoing momenta and $p_{3},$ $p_{4}$ are the outgoing momenta. In particular, $s$ is equal to the centre-of-mass energy squared, while $t$ is the momentum transfer squared. Since in this work we only focus on a gapless theory, we have $s+t+u=0.$ The scattering amplitude can be chosen to be a function of two Mandelstam variables only, e.g. $s$ and $t$, thus $A(s,t,u)=A(s,t)$.
2 Dispersion relations for exponentially bounded amplitudes
@@
The central object of our discussion is the $2\rightarrow 2$ scattering amplitude $A(s,t)$ in some gapless scalar theory. In particular, we assume that its low-energy behavior can be captured by the following EFT expansion:
@@
$+g_{5}(stu)(s^{2}+t^{2}+u^{2})+g_{6}(s^{2}+t^{2}+u^{2})^{3}+g^{\prime}_{6}(stu)^{2}+\cdots\,,$
which is manifestly crossing symmetric and vanishes at $(s,t)=(0,0)$; the couplings $g_{n}$ are known as Wilson coefficients.
Polynomially bounded amplitude.
In the past literature [3, 54], the Wilson coefficients $g_{n}$ have been constrained assuming that the UV completion is local, i.e. the amplitude is polynomially bounded in all complex directions in the $s$ and $t$ planes. This requirement together with partial-wave unitarity implies the Froissart-Martin bound in Eq. (1). In addition to these ingredients, the assumption of analyticity of the amplitude allows to derive various dispersion relations at fixed $t\leq 0$ connecting the Wilson coefficients in $A_{\rm EFT}$ with positive contributions coming from the discontinuities along the branch cuts of the amplitude. If we choose the contour $\gamma_{1}\cup\gamma_{2}$ in Fig. 1, we can write [6, 24]
@@
$\int_{C_{1}\cup C_{2}}\frac{{\rm d}\mu}{2\pi i}\frac{A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}=$ $\,(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}$ (5)
$+(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-u_{1})(\mu-u_{2})(\mu-u)}\,,$
@@
${\rm Disc}_{s}A(\mu,t)=\lim_{\epsilon\rightarrow 0^{+}}\frac{1}{2i}\left[A(\mu+i\epsilon,t)-A(\mu-i\epsilon,t)\right],$ (6)
and we have used crossing symmetry to write ${\rm Disc}_{s}A(-\mu-t,t)=-{\rm Disc}_{s}A(\mu,t)$ and defined analogue subtraction points in terms of the Mandelstam variable $u,$ i.e. $u_{1}=-s_{1}-t$ and $u_{2}=-s_{2}-t.$ The scale $\Lambda$ represents the cutoff scale of the EFT.
Figure 1: Integration contour in the complex $\mu$ plane used to derive the twice-subtracted dispersion relations in eqs. (7), (9) and (11). The initial contour $\gamma_{1}\cup\gamma_{2}$ on the left is taken of infinite radius. It is then deformed into the contour on the right, where the circle $C_{1}\cup C_{2}$ is centered at $\mu=-t/2.$ The boxes denote the two branch points of the $s$-channel (from $\mu=0$ to +$\infty$+ ∞) and $u$-channel (from $\mu=-t$ to $\infty$- ∞) branch cuts, respectively. The point $s$, $s_{1}$ and $s_{2}$ are the poles for which the residue theorem is used to derive the dispersion relations.
By computing the integral on the left-hand-side of Eq. (5) through the residue theorem, we obtain the twice-subtracted dispersion relation
$A_{\rm EFT}(s,t)=$ $\,(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}$ (7)
$+(s-s_{1})(s-s_{2})\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-u_{1})(\mu-u_{2})(\mu-u)}+a(t)+sb(t)\,,$
where
$a(t)\equiv\frac{1}{s_{1}-s_{2}}\left[s_{1}A_{\rm EFT}(s_{2},t)-s_{2}A_{\rm EFT}(s_{1},t)\right]\,,\qquad b(t)\equiv\frac{1}{s_{1}-s_{2}}\left[A_{\rm EFT}(s_{1},t)-A_{\rm EFT}(s_{2},t)\right]$ (8)
@@
are the subtraction terms. Note that we wrote $A_{\rm EFT}(s,t)$ because the left-hand-side of Eq. (7) satisfies the equation only for $s<\Lambda^{2}.$ This aspect is what allows us to make the connection with the low-energy EFT expansion (3). In fact, more precisely, the left-hand-side of Eq. (7) reads $A(s,t)|_{s<\Lambda^{2}}\simeq A_{\rm EFT}(s,t).$
To work with simpler formulas, in what follows we choose that the two subtraction points coincide and are equal to zero, that is, $s_{1}=s_{2}=0.$ In this way, the dispersion relation (7) becomes
where now the subtraction terms read
$a(t)=A_{\rm EFT}(0,t)\,,\qquad b(t)=\left.\frac{\partial A_{\rm EFT}(s,t)}{\partial s}\right|_{s=0}\,.$ (10)
Taking the second and higher order derivatives at $s=0$ of the dispersion relation (7) or (9), and combining the expressions into a set of integral inequalities, one can derive constraints on $g_{n}$ [54].
Exponentially bounded amplitude.
@@
In this work, we are interested in non-local UV completions in which the amplitude does not satisfy the property of polynomial boundedness. In particular, for fixed momentum transfer we assume that its growth for $|s|\rightarrow\infty$ is bounded as in Eq. (2). In this case the requirements of analyticity and partial-wave unitarity can still be satisfied, that is, the only pole singularities and branch cuts are confined to the real axis and the discontinuity across the branch cut is positive. However, in this case the integral over $\gamma_{1}\cup\gamma_{2}$ in Eq. (4) could blow up. [^3] It is worth to mention that for strictly localizable QFTs, in which the amplitude is bounded as (7) with $g<1/4,$ it can actually be shown that the amplitude cannot grow faster than a polynomial in the complex $s$ [67]. In this paper, instead, we are interested in non-localizable QFTs with $g>1/4$ (in particular $g=1$), for which a similar result does not necessarily apply. Therefore, modified dispersion relations that take into account a non-localizable exponential growth are needed. This means that we need to implement a modified dispersion relation. Let us start considering the forward limit $t=0,$ and later on we will address the case of non-vanishing momentum transfer that will require a more careful treatment.
@@
When $t=0$, we can assume that the amplitude is bounded as $A(s,t=0)<s^{2}\exp{(-\alpha s^{2g})}$ for any $s\in\mathbb{C},$ up to subdominant factors. This means that along the real axis the amplitude is exponentially suppressed in the forward limit, and can grow exponentially along some direction in the complex $s$ plane. To write down a dispersion relation for this type of amplitude, we can adopt the following strategy. We multiply the amplitude by $\exp{(\alpha s^{2g})}$ and define $A^{\prime}(s,t)\equiv\exp{(\alpha s^{2g})}A(s,t)$, which will be bounded by $|s|^{2}$ in the whole complex $s$ plane. In other words, we compensate the essential singularity at infinity, so that we can write down dispersion relations for $A^{\prime}(s,t=0)$ that are analogue to Eq. (9). Indeed, we have
@@
where the subtraction terms $a^{\prime}(0)$ and $b^{\prime}(0)$ have the same form as those in eq. (10), but with $A$ replaced by $A^{\prime}$. However, since our amplitude satisfies $A_{\rm EFT}(0,0)=0$ and $\partial_{s}A_{\rm EFT}(s,0)|_{s=0}=0,$ the two subtraction terms actually vanish in the forward limit, i.e. $a^{\prime}(0)=0=b^{\prime}(0).$
The relation between the Wilson coefficients in $A_{\rm EFT}$ and the positive moments – i.e. the derivatives with respect to $s$ evaluated at $s=0$ of the amplitude – will now be modified due to additional $\alpha$-dependent terms. In the next section we show how this affects the resulting EFT constraints.
We will also provide explicit examples of amplitudes that decay exponentially in the UV along physical directions and, at the same time, satisfy partial-wave unitarity, causality and crossing symmetry. For instance, the expression
$\begin{split}A(s,t,u)\sim\,C&\left[e^{-\alpha s^{2}}\log{(\sqrt{\alpha}s)}\log{(-\sqrt{\alpha}s)}+e^{-\alpha t^{2}}\log{(\sqrt{\alpha}t)}\log{(-\sqrt{\alpha}t)}+\right.\\ &\left.+e^{-\alpha u^{2}}\log{(\sqrt{\alpha}u)}\log{(-\sqrt{\alpha}u)}\right]\end{split}$ (12)
represents a possible UV behavior that is consistent with partial-wave unitarity for $s\gtrsim 1/\sqrt{\alpha}$ and that could be matched to the logarithmic running at low energies. In App. A we examine further examples, including amplitudes that are finite at the point $(s,t)=(0,0),$ which are therefore consistent with our working assumptions.
3 Positivity bounds at $t=0$
Let us now derive the positive bounds in the forward limit. Hereafter we study the non-localizable case with $g=1$. Taking two derivatives of the dispersion relation (11) with respect to $s$ and evaluating them at $s=0$, we obtain
$g_{2}=\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\rm{Disc}_{s}A^{\prime}(\mu,0)\frac{1}{\mu^{3}}\,.$ (13)
Unitarity implies the positivity of the last integral, thus we obtain the first constraint on the lowest Wilson coefficient, i.e. $g_{2}>0$, which remains the same as in the standard local case.
If we define
@@ 3 Positivity bounds at $t=0$
and analogue relations can be obtained for higher-order Wilson coefficients.
@@
Working directly with the last relations is not convenient due to the presence of negative contributions on the right-hand-sides. However, we can instead manipulate those equations and work with the following positive combinations of Wilson coefficients,
[…]
$\frac{1}{\mu^{9}}\right>=\,g_{8}+4\alpha g_{6}+\alpha^{2}g_{4}+\frac{\alpha^{3}}{6}g_{2}\,.$
[…]
$0<2\tilde{g}_{4}+\tilde{\alpha}<1,\qquad 0<4\tilde{g}_{6}+2\tilde{\alpha}\tilde{g}_{4}+\frac{1}{2}\tilde{\alpha}^{2}<2\tilde{g}_{4}+\tilde{\alpha},\qquad 4\tilde{g}_{6}+2\tilde{\alpha}\tilde{g}_{4}+\frac{1}{2}\tilde{\alpha}^{2}\geq(2\tilde{g}_{4}+\tilde{\alpha})^{2},$ (17)
where we have defined the dimensionless quantities $\tilde{\alpha}\equiv\alpha\Lambda^{4},$ $\tilde{g}_{4}\equiv g_{4}\Lambda^{4}/g_{2}$, $\tilde{g}_{6}\equiv g_{6}\Lambda^{8}/g_{2}$, and so on. In the $\alpha\rightarrow 0$ limit, these inequalities reduce to the well-known local case [3]. We also verified that using two Cauchy-Schwartz inequalities to derive the bounds on $\tilde{g}_{4}$ and $\tilde{g}_{6}$ is already sufficient to obtain the most optimal constraints, which is in agreement with other analyses performed in local cases [69].
The regions in the $\tilde{g}_{4}$-$\tilde{g}_{6}$ plane that are allowed by the constraints in (17) are shown in Fig. 2a for different values of $\tilde{\alpha}$. We can track how the allowed region changes as a function of $\tilde{\alpha}$, going from $\tilde{\alpha}=0$ to $\tilde{\alpha}=1.0$. In particular, we found that the regions corresponding to $\tilde{\alpha}\geq\sqrt{2}-1\approx 0.41$ have no overlap with the local case $\tilde{\alpha}=0$.
In Fig. 2b we plotted the allowed regions in the three-dimensional parameter space spanned by $\tilde{g}_{4}$, $\tilde{g}_{6}$ and $\tilde{g}_{8}$.
4 Bounds on $g_{3}$ coupling
Although the above derivation of positivity bounds in the forward limit is quite straightforward, it does not give us any information about the couplings in the EFT amplitude (3) that disappear when $t=0$. Constructing the dispersion relation for $t<0$ requires more details about the structure of the UV amplitude. For example, at finite $t$ it may be not possible to compensate the essential singularities in dispersion relation in a crossing-symmetric way. To avoid these potential obstacles, we assume a crossing-symmetric form of the amplitude which is consistent with partial-wave unitary and is inspired by the explicit examples provided in App. A. In particular, if the amplitude can be represented as a sum of $s$, $t$, and $u$ channels weighted with the corresponding exponentials [^4] Here we assume $F(s,t)$ to be bounded by $s^{2}$ in the entire complex plane, so that it can be expressed in terms of a twice-subtracted dispersion relation.,
$A(s,t,u)=F(t,u)e^{-\alpha s^{2g}}+F(s,u)e^{-\alpha t^{2g}}+F(s,t)e^{-\alpha u^{2g}},$ (18)
we can derive its $s\leftrightarrow u$ crossing-symmetric representation (specializing to $g=1$),

### v2

@@
fDepartment of Physics, Blackett Laboratory, Imperial College London, SW7 2AZ London, UK
The requirements of unitarity and causality lead to significant constraints on the Wilson coefficients of an EFT expansion, known as positivity bounds. Their standard derivation relies on the crucial assumption of polynomial boundedness on the growth of scattering amplitudes in the complex energy plane, which is a property satisfied by local QFTs, and by weakly coupled string theory in the Regge regime. The scope of this work is to clarify the role of locality by deriving generalized positivity bounds under the assumption of exponential boundedness, typical of non-local QFTs where the Froissart-Martin bound is usually not satisfied. Using appropriately modified dispersion relations, we derive new constraints and find regions in the EFT parameter space that do not admit a local UV completion. Furthermore, we show that there exist EFTs that satisfy IR causality and at the same time can admit a non-local UV completion, provided that the energy scale of non-locality is of the same order or larger than the EFT cutoff. Finally, we provide an explicit example of an exponentially bounded amplitude that satisfies partial-wave unitarity and asymptotic causality.
@@
Introduction.— To constrain the parameter space of viable effective field theories (EFTs) it is often not necessary to know the precise form of the UV completion. In many cases it is sufficient to assume certain physical properties in the UV regime, which can be translated into mathematical requirements on the $S$-matrix, from which restrictions on low-energy physics can then be derived [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] (see also [54] for a recent review). For example, standard requirements that are usually assumed are those of Lorentz invariance, unitarity and causality. These properties are encoded in a set of mathematical conditions, such as positivity of the imaginary part of the partial waves and analyticity of the $2\rightarrow 2$ scattering amplitude in the complex energy plane [55, 56, 57, 58].
A further assumption that is typically made on the UV completion is that of locality, which states that the growth of the scattering amplitude is polynomially bounded in the complex energy plane. For local gapped theories it can be proven that the amplitude for fixed physical momentum transfer is polynomially bounded by $|s|^{2}$ in the whole complex plane; this is known as Froissart-Martin bound [^1] It is worth to mention that Froissart [59] and Martin [60] initially showed that the amplitude was bounded by $|s|$ in the forward limit $t=0.$ Subsequently, Jin and Martin [61] showed that the amplitude for generic physical momentum transfer $t\leq 0$ is bounded by $|s|^{2}.$ In any case, in this work we refer to the latter as Froissart-Martin bound. [59, 60, 61]. Although for theories with massless states no rigorous general proof of this bound exists, it is common to assume that the growth of the amplitude for fixed physical $t$ is bounded by $|s|^{2}$ in the whole complex plane: [^2] This can be verified at least for graviton-mediated amplitudes in $d>4$, using the method of eikonal resummation [62, 63].
@@
where $B$ is a subdominant factor that may depend on $s$.
To clarify the role of locality, it is instructive to investigate what happens when we allow for a non-polynomial growth of the scattering amplitude, i.e. for violations of the Froissart-Martin bound (1). In particular, we are going to consider amplitudes that are exponentially bounded in the complex $s$ plane for fixed physical $t$:
@@
where $\alpha$ and $g$ are two positive constant parameters, and $C$ a subdominant factor that may depend on $s$.
@@
QFTs in which scattering amplitudes are characterized by this type of exponential behavior could be classified according to Jaffe’s criterion [64, 65, 66, 67, 68]: if $g<1/4$ the corresponding QFT is said to be localizable, whereas if $g\geq 1/4$ the QFT is said to be non-localizable. When $g=0$ we recover the standard local QFT description with polynomially bounded amplitudes. The name non-localizability is related to the fact that the fields must be delocalized through smeared test functions in order to make the position-space correlators (such as the Wightman function) finite and well-defined. In what follows we use the words local (non-local) and localizable (non-localizable) interchangeably.
The aim of this Letter is to understand what type of constraints can be derived on the Wilson coefficients of the low-energy EFT expansion if the UV completion is assumed to be non-localizable according to (2), and study their physical implications. We work in four spacetime dimensions, and adopt the mostly plus metric signature ($+$)(-$+$)( - + + + ) and the Natural units system $(\hbar=1=c).$
@@
Dispersion relations.— The central object of our investigation is the elastic $2\rightarrow 2$ scattering amplitude $A(s,t,u)$ in some gapless scalar theory, where $s=-(p_{1}+p_{2})^{2},$ $t=-(p_{1}-p_{3})^{2},$ $u=-(p_{1}-p_{4})^{2}$ are the Mandelstam variables satisfying the relation $s+t+u=0$; $p_{1},$ $p_{2}$ are the ingoing momenta and $p_{3},$ $p_{4}$ are the outgoing momenta. In particular, $s$ is the centre-of-mass energy squared, while $t$ is the momentum transfer squared. The scattering amplitude can be chosen to be a function of two Mandelstam variables only, e.g. $A(s,t,u)=A(s,t)$.
@@
We assume that the low-energy behavior of the amplitude can be captured by the following EFT expansion:
@@
$+g_{5}(stu)(s^{2}+t^{2}+u^{2})+g_{6}(s^{2}+t^{2}+u^{2})^{3}+\cdots\,,$
which is manifestly crossing symmetric; the couplings $g_{n}$ are called Wilson coefficients.
In the past literature [3, 54], the Wilson coefficients have been constrained assuming that the UV completion is local, i.e. the amplitude is polynomially bounded in all complex directions in the $s$ and $t$ planes. In addition to these ingredients, the assumption of analyticity of the amplitude allows to derive various dispersion relations at fixed $t\leq 0$, connecting the Wilson coefficients in $A_{\rm EFT}$ with positive contributions coming from the discontinuities along the branch cuts of the amplitude. Indeed, by choosing the contour $\gamma_{1}\cup\gamma_{2}$ in Fig. 1, we can write [6, 24]
@@
$\int_{C_{1}\cup C_{2}}\frac{{\rm d}\mu}{2\pi i}\frac{A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}=$ $\,\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-s_{1})(\mu-s_{2})(\mu-s)}$ (5)
$+\int_{\Lambda^{2}}^{\infty}\frac{{\rm d}\mu}{\pi}\frac{{\rm Disc}_{s}A(\mu,t)}{(\mu-u_{1})(\mu-u_{2})(\mu-u)}\,,$
where ${\rm Disc}_{s}A(\mu,t)=\lim_{\epsilon\rightarrow 0^{+}}\frac{1}{2i}[A(\mu+i\epsilon,t)-A(\mu-i\epsilon,t)];$ we have used crossing symmetry to write ${\rm Disc}_{s}A(-\mu-t,t)=-{\rm Disc}_{s}A(\mu,t)$ and have defined the similar subtraction points in terms of the Mandelstam variable $u,$ i.e. $u_{1}=-s_{1}-t$ and $u_{2}=-s_{2}-t.$ The scale $\Lambda$ represents the EFT cutoff.
Figure 1: Integration contour in the complex $\mu$ plane used to derive the twice-subtracted dispersion relations in Eqs. (6) and (8) (with $s_{1}=s_{2}=0$). The initial contour $\gamma_{1}\cup\gamma_{2}$ on the left is taken of infinite radius. It is then deformed into the contour on the right, where the circle $C_{1}\cup C_{2}$ is centered at $\mu=-t/2.$ The boxes denote the two branch points of the $s$-channel (from $\mu=0$ to +$\infty$+ ∞) and $u$-channel (from $\mu=-t$ to $\infty$- ∞) branch cuts, respectively. The point $s$, $s_{1}$ and $s_{2}$ are the poles in Eq. (4).
By computing the integral on the left-hand-side of Eq. (5) through the residue theorem, we can obtain the twice-subtracted dispersion relation. For simplicity we choose $s_{1}=s_{2}=0,$ thus we can write
@@
$a(t)=A_{\rm EFT}(0,t)\,,\qquad b(t)=\left.\frac{\partial A_{\rm EFT}(s,t)}{\partial s}\right|_{s=0}\,,$ (7)
@@
are the two subtraction terms. Note that, we wrote $A_{\rm EFT}(s,t)$ because the left-hand-side of Eq. (6) satisfies the equation only for $s<\Lambda^{2}.$ This aspect is what allows us to make the connection with the low-energy EFT expansion (3). In fact, more precisely, the left-hand-side of Eq. (6) should read $A(s,t)|_{s<\Lambda^{2}}\simeq A_{\rm EFT}(s,t).$
@@
If the assumption of polynomial boundedness is given up, the integral over $\gamma_{1}\cup\gamma_{2}$ in Eq. (4) could blow up. [^3] It is worth to mention that for localizable QFTs, in which the amplitude is bounded as (2) with $g<1/4,$ it can actually be shown that the amplitude cannot grow faster than a polynomial in the complex $s$ plane [67]. In this Letter, instead, we are interested in non-localizable QFTs with $g>1/4$ (in particular $g=1$), for which a similar result does not necessarily apply. Therefore, modified dispersion relations that take into account a non-localizable exponential growth are needed in our case. This means that we need to implement a modified dispersion relation. Let us start considering the forward limit $t=0,$ and later on we will address the case of non-vanishing physical momentum transfer that will require a more careful treatment.
@@
When $t=0$, we can assume that the amplitude is bounded as $A(s,t=0)<s^{2}\exp{(-\alpha s^{2g})}$ for any $s\in\mathbb{C},$ up to subdominant factors. This means that along the real axis the amplitude is exponentially suppressed in the forward limit, and can grow exponentially along some direction in the complex $s$ plane. To write down a dispersion relation for this type of amplitude, we can adopt the following strategy. We multiply the amplitude by $\exp{(\alpha s^{2g})}$ and define $A^{\prime}(s,t)\equiv\exp{(\alpha s^{2g})}A(s,t)$, which will be bounded by $|s|^{2}$ in the whole complex $s$ plane. In other words, we compensate for the essential singularity at infinity, so that we can write a dispersion relation for $A^{\prime}(s,t=0)$ that is analogous to Eq. (6):
@@
where $a^{\prime}(0)$ and $b^{\prime}(0)$ have the same form as those in Eq. (7), but with $A$ replaced by $A^{\prime}$. However, since our amplitude satisfies $A_{\rm EFT}(0,0)=0$ and $\partial_{s}A_{\rm EFT}(s,0)|_{s=0}=0,$ the two subtraction terms actually vanish in the forward limit, i.e. $a^{\prime}(0)=0=b^{\prime}(0).$
Positivity bounds for $\boldsymbol{t=0}$.— Let us now derive the positivity bounds in the forward limit. Hereafter we study the non-localizable case with $g=1$.
If we define the positive quantity
@@
$0<2\tilde{g}_{4}+\tilde{\alpha}$ $<1,\qquad 0<4\tilde{g}_{6}+2\tilde{\alpha}\tilde{g}_{4}+\tilde{\alpha}^{2}/2<2\tilde{g}_{4}+\tilde{\alpha}\,,$ (11)
$4\tilde{g}_{6}+2\tilde{\alpha}\tilde{g}_{4}+\tilde{\alpha}^{2}/2\geq(2\tilde{g}_{4}+\tilde{\alpha})^{2}\,,$
where we have defined the dimensionless quantities $\tilde{\alpha}\equiv\alpha\Lambda^{4},$ $\tilde{g}_{4}\equiv g_{4}\Lambda^{4}/g_{2}$, $\tilde{g}_{6}\equiv g_{6}\Lambda^{8}/g_{2}$, and so on. In the $\alpha\rightarrow 0$ limit, we recover the well-known local case [7].
@@
The regions in the $\tilde{g}_{4}$-$\tilde{g}_{6}$ plane that are allowed by the constraints in (11) are shown in Fig. 2a for different values of $\tilde{\alpha}$. We can track how the allowed region changes as a function of $\tilde{\alpha}$, going from $\tilde{\alpha}=0$ to $\tilde{\alpha}=1.0$. In particular, we found that the regions corresponding to $\tilde{\alpha}\geq\sqrt{2}-1\approx 0.41$ have no overlap with the local case $\tilde{\alpha}=0$. We also verified that using two Cauchy-Schwartz inequalities to derive the bounds on $\tilde{g}_{4}$ and $\tilde{g}_{6}$ is already sufficient to obtain the most optimal constraints, in agreement with the analyses performed in local cases [7, 69]. In Fig. 2b we plotted the allowed regions in the three-dimensional parameter space spanned by $\tilde{g}_{4}$, $\tilde{g}_{6}$, $\tilde{g}_{8}$.
Positivity bounds for $\boldsymbol{t<0}$.— The above derivation of positivity bounds does not give us any information about the couplings in the EFT amplitude (3) that disappear when $t=0$, such as $g_{3}$. Constructing the dispersion relation for $t<0$ requires more details about the structure of the UV amplitude. We assume the following crossing-symmetric form, which is consistent with partial-wave unitary (see the Supplemental Material):
$A(s,t,u)=F(t,u)e^{-\alpha s^{2}}+F(s,u)e^{-\alpha t^{2}}+F(s,t)e^{-\alpha u^{2}}\,,$ (12)
where $F(s,t)$ is bounded by $s^{2}$ in the entire complex plane. We can then write the $s\leftrightarrow u$ crossing-symmetric representation as